OpenStudy (anonymous):

I have folded one of the pages randomly of a book . The sum of the page numbers except the page which is folded is 15000. what are the pages that are folded?

OpenStudy (anonymous):

What ? Can you answer better ?!!!

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (akashdeepdeb):

Argghhh.. I am getting one equation at this......

OpenStudy (anonymous):

what is it

OpenStudy (akashdeepdeb):

If one page is folded that means x and the previous page of the same leaf is x-1 Forming an arithmetic progression for sum of consecutive numbers: n(n+1)/2 = 15000 + x + x-1 Understand this? :)

OpenStudy (anonymous):

ya but how would u solve it

OpenStudy (akashdeepdeb):

Yes..That's what has been boggling my mind for some hours now... Is there anything else given in the question? We need to get another equation! :D

OpenStudy (anonymous):

sorry but nothing else was given in it but I am knowing the ans its 25 & 26

OpenStudy (akashdeepdeb):

Tag some people see if they can help you or not! :D Tag them by putting '@'*theirusername* :)

OpenStudy (anonymous):

i dont know how to tag someone

OpenStudy (akashdeepdeb):

@Ayushi2103 ^ LIKE THAT! :D

OpenStudy (anonymous):

but where to type it

OpenStudy (anonymous):

ok i got it thanx

OpenStudy (anonymous):

@goformit100

OpenStudy (akashdeepdeb):

Try this: 15000 = (n-2)/2 * [a + (n-2 -1)d] where a=1,d=1 :D

OpenStudy (anonymous):

does this work

OpenStudy (akashdeepdeb):

I think so.. Because the sum is 15000 without including 2 pages so n-2 :D

OpenStudy (anonymous):

i got the solution n(n+1)/2=>15000 by hit and trial 1+2+3...+173=15051 x+x+1=15051-15000 2x+1=51 2x=50 x=25 x+1=26

OpenStudy (anonymous):

thanx for giving me the clue and helping me with it

OpenStudy (akashdeepdeb):

Oh Awesome! :D

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

we ll meet later on cauz i wanna go bye